OpenStudy (anonymous):

How would I solve this for x? 12/42 = x/14

OpenStudy (anonymous):

x=4, sorry

OpenStudy (johnweldon1993):

You would multiply both sides of the equation by 14.....when you multiply say 11/14 by 14...the 14's cancel and all you have is 11 So for you (12/42) * 14 = (x/14) * 14 you're left with (12/42) *14 = x

OpenStudy (johnweldon1993):

Which yes...will come out to @kutabs answer of 4 Understandable??

OpenStudy (anonymous):

Why am I multiplying both by 14?

OpenStudy (johnweldon1993):

like I said....you want to isolate x.....and to do that....since x is being DIVIDED by 14 here....you do the INVERSE which is MULTIPLYING by 14....this cancels the dividing out and leaves you with JUST x on that right side

OpenStudy (anonymous):

because you need the value of x, and you've got the value of x/14 (equal to 12/42). Now, to get the value of x multiply both sides by 14 to that the RHS becomes (x/14)*14= x

OpenStudy (anonymous):

Oh - I see. Danke!

OpenStudy (anonymous):

@johnweldon1993 - http://openstudy.com/study#/updates/519fae96e4b0e9c70c32e3e0 If you can solve this please. I am not being able to, and it's pissing me off!